The base of a 20 foot ladder is placed 8 feet from a wall.How far up the wall will the ladder reach?

Answer by checkley79(3341)

You have a right triangle.
The hypotenuse is the length of the ladder (c).
T height of the ladder is b.
a^2+b^2=c^2
8^2+b^2=20^2
64+b^2=400
b^2=400-64
b^2=336
b=sqrt336
b=18.33 ft. ans.